A solution of Ba(OH)2 was standardized against 0.1016 g of primary standard 

benzoic acid C6H5COOH. Calculate the concentration of Ba(OH)2 if the end point 

was 20.89 mL. 


1
Expert's answer
2021-08-11T05:23:03-0400

Solution:

The titration reaction is given by the following chemical equation:

Ba(OH)2 + 2C6H5COOH → Ba(C6H5COO)2 + 2H2O

According to the equation above:

Moles of Ba(OH)2 = (Moles of C6H5COOH) / 2


The molar mass of C6H5COOH is 122.12 g/mol.

Therefore,

(0.1016 g C6H5COOH) × (1 mol C6H5COOH / 122.12 g C6H5COOH) = 0.000832 mol C6H5COOH


Moles of Ba(OH)2 = (Moles of C6H5COOH) / 2 = (0.000832 mol) / 2 = 0.000416 mol Ba(OH)2


Molarity = Moles / Solution volume

Molarity of Ba(OH)2 = (0.000416 mol) / (20.89×10-3 L) = 0.0199 mol/L = 0.0199 M

Molarity of Ba(OH)2 = 0.0199 M


Answer: The concentration of Ba(OH)2 is 0.0199 mol/L